OpenStudy (anonymous):

doing the ratio test and trying to following pauls online notes.com and confused about how 4^(2n+1)/4^(2n+3) simplifies down to 4^2....am i missing something basic? i know 4^n+3 / 4^n+2 is 4^n...

hartnn (hartnn):

4^(2n+1)/4^(2n+3) = 4^(2n+1 - 2n-3) = 4^(-2) = 1/16

hartnn (hartnn):

the ratio test ? so you will do \(a_{n+1}/a_n\) right >

OpenStudy (anonymous):

oh. ok. thanks. i was missing something simple. and yes, that part of the ratio test i understand. i just sometimes have a hard time seeing certain things. i'm borderline learning disabled so thanks for spelling it out for me. that helps.